Mr. Malone wants to change the overall risk of his portfolio. Currently, his portfolio is a...

  1. Mr. Malone wants to change the overall risk of his portfolio. Currently, his portfolio is a combination of risky assets with a beta of 1.25 and an expected return of 14%. He will add a risk-free asset (U.S. Treasury bill) to his portfolio. If he wants a beta of 1.0, what percent of his wealth should be in the risky portfolio and what percentage should be in the risk-free asset? If he wants a beta of 0.75? If he wants a beta of 0.50? If he wants a beta of 0.25? Is there a pattern here?

PLEASE GIVE ME A FULL EXPLANATION WITH ALL THE FORMULAS, THANK YOU
Answer #1

Answer-

Existion Beta = 1.25, Expected return= 14%

1.Calculation of risk free asset to be purchased to reduce existing beta to 1.00

Let assume that weight of portfolio be 'w' and weight of risk free asset be '1-w'

Hence,

portfolio beta= weighted avg. beta

1.00 = w*1.25+(1-w)*0

1.00 = 1.25w+0

1.00 = 1.25w

w = 1.00/1.25

weight of risky asset = 0.80

weight of risk free asset = 1-0.80

= 0.20

2. Calculation of risk free asset to be purchased to reduce existing beta to 0.75

Let assume that weight of portfolio be 'w' and weight of risk free asset be '1-w'

Hence,

portfolio beta= weighted avg. beta

0.75 = w*1.25+(1-w)*0

0.75 = 1.25w+0

0.75 = 1.25w

w = 0.75/1.25

weight of risky asset = 0.60

weight of risk free asset = 1-0.60

= 0.40

3.Calculation of risk free asset to be purchased to reduce existing beta to 0.50

Let assume that weight of portfolio be 'w' and weight of risk free asset be '1-w'

Hence,

portfolio beta= weighted avg. beta

0.50 = w*1.25+(1-w)*0

0.50 = 1.25w+0

0.50 = 1.25w

w = 0.50/1.25

weight of risky asset = 0.40

weight of risk free asset = 1 - 0.40

= 0.60

4.Calculation of risk free asset to be purchased to reduce existing beta to 0.25

Let assume that weight of portfolio be 'w' and weight of risk free asset be '1-w'

Hence,

portfolio beta= weighted avg. beta

0.25 = w*1.25+(1-w)*0

0.25 = 1.25w+0

0.25 = 1.25w

w = 0.25/1.25

weight of risky asset = 0.20

weight of risk free asset = 1 - 0.20

= 0.80

SUMMARY

Sr. No Expected Beta

Weight of Risk

free asset

Weight of Risky

asset

1 1.00 0.20 0.80
2 0.75 0.40 0.60
3 0.50 0.60 0.40
4 0.20 0.80 0.20